# Improving LLM Agents with Reinforcement Learning on Cryptographic CTF Challenges

### v1

@@ Abstract
Large Language Models (LLMs) still struggle with the structured reasoning and tool-assisted computation needed for problem solving in cybersecurity applications. In this work, we introduce random-crypto, a cryptographic Capture-the-Flag (CTF) challenge generator framework that we use to fine-tune a tool-augmented Llama-3.1-8B with Guided Reinforcement Prompt Optimisation (GRPO), allowing the agent to iteratively write and execute Python inside an isolated REPL. GRPO yields a +53% absolute jump in Pass@8 on unseen random-crypto tasks (0.35 → 0.88) and raises Majority@8 to 0.41. The fine-tuned agent also generalizes to an external dataset. On a subset of picoCTF cryptography problems, it improves Pass@8 by +13 pp. Ablations show the gains stem from more reliable tool invocation and code synthesis, rather than superficial prompt adaptation. We released the benchmark and training code on GitHub. [^1] https://github.com/aielte-research/HackSynth-GRPO
@@ 1 Introduction
Large language models (LLMs) have rapidly evolved from powerful text generators to sophisticated agents capable of performing complex reasoning tasks that frequently surpass human baselines [27]. This rapid advancement has spurred significant research interest in deploying LLM-based agents in critical cybersecurity roles, such as autonomous penetration testing and vulnerability detection [7, 6, 31].
Penetration testing has garnered considerable attention due to its practical implications for system security. This approach involves simulating cyberattacks to identify vulnerabilities and subsequently documenting exploitable weaknesses. Current LLM-based penetration testing agents primarily emphasize autonomous hacking capabilities, with their effectiveness typically assessed using Capture The Flag (CTF) challenges—structured cybersecurity puzzles designed to emulate real-world offensive and defensive scenarios [28, 15, 24, 30]. However, existing benchmarks remain limited in scale and diversity [24], posing a challenge for rigorous evaluation and comprehensive training of LLM-based cybersecurity agents.
Reinforcement learning (RL) has emerged as a promising approach to enhance the capabilities of LLM agents, particularly through structured fine-tuning methods that can improve multi-step reasoning and decision-making abilities [2]. The Guided Reinforcement Prompt Optimization (GRPO) [25] method introduced by the DeepSeek team has notably demonstrated state-of-the-art performance improvements in structured reasoning tasks while maintaining computational efficiency [9].
We introduce the random-crypto benchmark—a procedurally generated dataset comprising diverse cryptographic CTF challenges. This dataset provides an effectively arbitrarily large training corpus, facilitating the systematic exploration of RL-based improvements. In this study, we use the random-crypto dataset to fine-tune a tool-augmented Llama-3.1-8B model [13]. We demonstrate that GRPO, not only enhances agent performance in cryptographic problem-solving but also improves the reliability and effectiveness of tool usage.
Specifically, our paper makes the following contributions:
@@
    We introduce a novel, programmatically generated cryptographic benchmark (random-crypto) for training and evaluation of cybersecurity-focused language models.
@@
    We demonstrate performance enhancements in LLM agents through GRPO-based reinforcement learning, particularly highlighting improvements in tool-augmented interactions.
@@
    We validate the generalization of our trained agents on an external dataset (picoCTF), underscoring the robustness of our approach.
@@ 2.1 CTF Challenges
Cryptographic challenges in particular revolve around exploiting weaknesses or logical flaws in encryption and encoding methods. These challenges commonly involve classical ciphers (e.g., Caesar, Vigenère, substitution), symmetric-key encryption (e.g., AES, DES), and public-key encryption schemes (e.g., RSA, ECC). Participants may also encounter tasks involving hashing algorithms, digital signatures, and custom cryptosystems, which often require creative approaches and deep theoretical understanding to solve.
@@ 2.2 LLM Agents in Cybersecurity
Recent advancements in large language models (LLMs) have sparked significant interest in their application within cybersecurity, particularly in penetration testing and security assessments. Researchers have adapted Capture The Flag (CTF) challenges as standardized benchmarks [28, 15, 30, 24], to systematically evaluate the cybersecurity capabilities of LLM-based agents. Initial evaluations revealed that general-purpose LLMs struggled to effectively handle specialized cybersecurity tasks due to their inherent inability to interact directly with computing environments [26, 28, 15, 4].
In other domains, to bridge this interaction gap, newer architectures introduced tool augmentation, empowering LLM agents to invoke external tools or APIs. Frameworks like ReAct (Reason+Act) demonstrated substantial improvements by integrating logical reasoning with actionable tool use [29]. OpenAI’s ChatGPT further extended this approach through a plugin/function-calling API, enabling deterministic JSON tool calls [16]. Additionally, Anthropic’s Model Context Protocol (MCP) provides a structured interface for models to interact seamlessly with external data sources and tools, facilitating complex cybersecurity tasks that require real-time computational support [1, 11]. These tool-augmented LLM agents have markedly advanced the state-of-the-art. However, they have yet to be fully utilized in autonomous penetration testing.
@@ 2.3 Reinforcement Learning to Enhance LLM Reasoning
Recent work has shown that traditional RLHF, while effective in improving helpfulness and safety, often falls short in domains that require structured, multi-hop reasoning like math, programming, and cybersecurity. To address this, specialized algorithms have emerged, such as Group Relative Policy Optimization (GRPO) [25]. GRPO is a variant of PPO tailored for reasoning-intensive tasks: instead of relying on scalar preference labels, it compares multiple candidate outputs and assigns relative rewards within a group. This relative evaluation strategy encourages diversity and robustness in generated solutions, helping the model refine its internal decision-making process. GRPO has been successfully applied to train DeepSeek-R1 [9], achieving state-of-the-art performance on mathematical reasoning and competitive coding benchmarks. Its low computational overhead and compatibility with low-rank adaptation methods such as QLoRA [5] make it an attractive choice for resource-constrained environments.
2.4 PicoCTF Evaluation Subset
To evaluate the transferability of the trained models beyond random-crypto, we curated a subset of the picoCTF benchmark [15]. Challenges were selected based on two criteria: (i) the challenge is tagged as cryptography, and (ii) the challenge provides a self-contained input, either as plaintext or as files that can be concatenated into an ASCII string. This excludes tasks requiring interactive services (e.g., netcat) or non-ASCII artifacts (e.g., images). Using these filters, we obtained 8 cryptography challenges for evaluation.
@@ 3.1 Random-Crypto Benchmark Dataset
We introduce the random-crypto benchmark, a systematically created, human-verified dataset tailored for training and evaluating cybersecurity-focused language models. The benchmark consists of cryptographic Capture The Flag (CTF) challenges designed to test a solver’s ability to exploit vulnerabilities in cryptographic schemes and protocols. Each challenge requires recovering a hidden flag in the form flag{...} by leveraging specific weaknesses.
@@
  • Hard: Complex tasks requiring advanced cryptographic knowledge, scripting capabilities, and potentially external online tools or frequency analysis methods (e.g., substitution cipher with frequency analysis).
@@
Each challenge is generated by first selecting a sub-type and then randomizing parameters specific to the cryptographic scheme, such as cipher keys or substitution dictionaries. Subsequently, a random flag is encrypted using the chosen parameters. An LLM-generated narrative accompanies each ciphertext, providing context and essential information to ensure solvability (see Appendix 0.B, Prompt 1). For example, an RSA challenge includes both the ciphertext and the public key in its narrative.
Manual validation was conducted by the authors, who solved at least one instance per sub-type to verify solvability. Parameter spaces were also reviewed to avoid trivial or unsolvable configurations. The dataset comprises 50 manually validated challenges used exclusively for testing, along with 5,000 automatically generated but unvalidated challenges designated for agent training.
@@ 3.2 Reinforcement Learning
We fine-tuned the Llama-3.1-8B [13] model on 1700 challenges from the easy difficulty subset using the Guided Reinforcement Prompt Optimization (GRPO) algorithm. The training specifically leveraged a structured interaction with a Python execution server via Anthropic’s Model Context Protocol (MCP), allowing the model up to four sequential tool invocations per challenge. The decision to restrict training to easy challenges was motivated by initial experiments showing that more difficult challenges negatively impacted the training process, causing the model to converge prematurely to local optima by prioritizing superficial reward types, thus reducing its overall problem-solving efficacy.
During each training step, we generated eight candidate trajectories for four distinct challenges, resulting in an effective batch size of 32. Each candidate trajectory consisted of iterative reasoning followed by structured JSON-formatted tool calls. The model interaction cycle included:
@@
The model’s learning process was guided by a composite reward function (see Table 1) designed to encourage correct and efficient problem-solving behavior. To address the challenge of the model prematurely optimizing towards superficial rewards (such as formatting or tool-calling correctness without genuine problem-solving), we implemented a deduction strategy. Specifically, we penalized scenarios where the model prematurely produced an answer immediately following a tool call, indicating hallucinated rather than computed outputs.
@@
Tool Calling Format Reward The model produces a valid JSON tool call 0.2
Python Execution Reward The MCP server executes the code without error 0.3
Training spanned $250$ steps and was facilitated by an Unsloth [10, 3] wrapper that managed multiple interactions within a single inference pass. The total token generation was limited to 8192 per interaction sequence, and training was augmented using QLoRA [5] adapters.
We explicitly instructed the model to strictly adhere to the interaction protocol via structured prompts (see Appendix 0.B, Prompt 2). The available Python execution server exposed three functionalities:
  • execute_python: Executes provided Python code and returns output.
  • list_variables: Lists variables currently stored in memory.
@@
4 Results
This section presents the evaluation of five leading LLMs: Llama-4-Scout-17B-16E [14], Llama-3.1-70B [12], Llama-3.1-8B [13], GPT-4.1 [17], and o3 [18]—on the randomized cryptographic benchmark. Additionally, we analyze the impact of reinforcement learning (RL)-based fine-tuning on model performance.
4.1 Benchmarking
We report two standard LLM evaluation metrics: Pass@8 is the proportion of tasks for which at least one out of eight independent model generations successfully solves the task, and Maj@8 is the proportion of tasks for which at least five out of eight independent model generations successfully solve the task.
Table 2 summarizes the experimental outcomes, measured by Pass@8 and Maj@8 metrics. GPT-4.1 and o3 models consistently outperform Llama-family models, highlighting the advanced reasoning capabilities of these architectures. Interestingly, supplementary hints generally improve performance for scenarios without tool usage, but this trend does not hold when tools are available. Particularly for smaller models like Llama-3.1-8B, the addition of hints adversely affects performance due to brute-force methods prompted by hints that cause the models to generate code that does not comply with environment constraints (memory and runtime). For larger and more capable models (GPT-4.1 and Scout), hints beneficially complement tool usage, leading to improved overall results. Differences in prompt engineering significantly affected the performance across models. For instance, the prompt was optimized for Llama-3.1-8B, inadvertently disadvantaging larger Llama models.
@@ 4.2 Reinforcement Learning
To investigate the impact of reinforcement learning (RL) on cryptographic task performance, we conducted three distinct training experiments using the Llama-3.1-8B model: (i) training without hints for 251 steps on easy challenges, (ii) training with hints for 251 steps, (iii) curriculum-based training, initially providing hints for 130 steps followed by 121 steps without hints.
At the outset, the model frequently produced valid Python code via tool calls but simultaneously hallucinated answers rather than executing the code. Consequently, even though the generated scripts could yield correct flags, the model’s preemptive hallucination terminated the interaction prematurely.
Figure 1 illustrates the progression of average rewards throughout training, showing an increase from approximately 0.6 to 0.9 across all scenarios. The hinted training regime consistently achieved higher rewards due to the hints guiding the model to the appropriate solution..
Metrics tracking successful Python code execution (defined as code running without errors) improved notably from roughly 39% to approximately 56%. The adherence to the correct answer format saw an even more substantial improvement, rising from about 40% to nearly 80%. However, properly formatted tool calls remained steady at approximately 95%, owing to extensive prompt engineering prior to training.
Figure 1: Reward gained during training.
Table 3: Pass@8 and Maj@8 results before and after GRPO training on the easy subset of random-crypto.
After Training Difference
@@
w/ Hints $0.88$ $0.41$ +0.53 +0.17
w/o Hints $0.41$ $0.24$ +0.06 +0.00
Curriculum $0.88$ $0.35$ +0.53 +0.11
Post-training results on the easy subset of the random-crypto benchmark (Table 3) reveal significant performance improvements. The curriculum and hinted models achieved the highest gains, indicating that training with hints effectively guides the model toward accurate solutions more frequently. Specifically, certain challenges that remained unsolved without hints were able to be solved when the model was equipped with hints, suggesting that hints effectively direct model focus toward solution paths.
Table 4 demonstrates that improvements gained from GRPO training generalized beyond the random-crypto dataset, as evidenced by enhanced performance on the picoCTF cryptographic benchmark [15]. Despite a limited set of picoCTF challenges constraining definitive conclusions, improved Python execution reliability positively impacted tasks external to the initial training data, implying genuine skill acquisition rather than dataset-specific overfitting.
Table 4: Pass@8 and Maj@8 results before and after GRPO training on PicoCTF cryptography challenges.
After Training Difference

### v2

@@ Abstract
We present Random-Crypto, a procedurally generated cryptographic Capture The Flag (CTF) dataset designed to unlock the potential of Reinforcement Learning (RL) for LLM-based agents in security-sensitive domains. Cryptographic reasoning offers an ideal RL testbed: it combines precise validation, structured multi-step inference, and reliance on reliable computational tool use. Leveraging these properties, we fine-tune a Python tool-augmented Llama-3.1-8B via Group Relative Policy Optimization (GRPO) in a secure execution environment. The resulting agent achieves a significant improvement in Pass@8 on previously unseen challenges. Moreover, the improvements generalize to two external benchmarks: picoCTF, spanning both crypto and non-crypto tasks, and AICrypto MCQ, a multiple-choice benchmark of 135 cryptography questions. Ablation studies attribute the gains to enhanced tool usage and procedural reasoning. These findings position Random-Crypto as a rich training ground for building intelligent, adaptable LLM agents capable of handling complex cybersecurity tasks.
keywords:
[…]
@@ 1 Introduction
Large Language Models (LLMs) are rapidly evolving from general-purpose text generators into autonomous agents—capable of complex reasoning and increasingly rivaling human performance across domains [1]. This rapid advancement has caused significant research interest in deploying LLM-based agents in critical cybersecurity contexts, such as autonomous penetration testing and vulnerability detection [2, 3, 4]. Penetration testing, which involves simulating cyberattacks to proactively identify and document exploitable vulnerabilities, is a valuable practice for enhancing system security. Current LLM-based penetration testing agents primarily focus on autonomous exploitation capabilities. Capture The Flag (CTF) challenges have emerged as a natural benchmark for these capabilities: they emulate real-world offensive and defensive scenarios, enforce precise success criteria, and demand structured multi-step reasoning [5, 6, 7, 8]. Despite advancements in scripted and tool-augmented approaches, a significant challenge remains: the absence of scalable, rigorous environments for training these agents, rather than just evaluating them.
Reinforcement Learning (RL) has emerged as a promising approach to enhance the capabilities of LLM agents and has shown strong results in mathematics, coding, and planning. This enhancement is particularly achieved through structured fine-tuning methods that improve multi-step reasoning and decision-making abilities [9]. The recently proposed Group Relative Policy Optimization (GRPO) [10], in particular, supports efficient fine-tuning using ranked outputs, and has shown strong performance on competitive reasoning tasks with a relatively modest compute effort [11]. However, practical RL training requires environments that provide abundant, objectively verifiable tasks, which are often lacking in cybersecurity datasets.
We address this gap with Random-Crypto, a procedurally generated dataset for cryptographic Capture The Flag (CTF) challenges that encompasses 50 different algorithmic families and over 5,000 unique tasks. Each challenge is randomized in its parameters, while embedding a randomized text flag, and is set within a contextual narrative generated by an LLM. This benchmark enables agent-centric, reward-driven training at scale, while preserving the symbolic accuracy of cryptographic problem-solving. This dataset provides an effectively limitless training corpus, allowing for extensive exploration and thorough evaluation of improvements based on RL. By focusing on cryptographic reasoning—a field that involves verifiable objectives, compositional tool use, and step-by-step logical inference—we create an environment where RL can lead agents not only to achieve correctness, but also to develop competence.
We utilize the Random-Crypto dataset to fine-tune a tool-augmented Llama-3.1-8B model [12], using a secure Python execution environment. Our results show that the enhanced agent significantly improves performance on previously unseen crypto tasks, increasing the Pass@8 metric from 0.35 to 0.88. Furthermore, the improvements generalize beyond the Random-Crypto distribution to two external benchmarks: picoCTF, which spans both cryptographic and non-cryptographic tasks, and AICrypto MCQ, a multiple-choice benchmark of 135 cryptography questions covering both theoretical and computational problems. In both cases, the model demonstrates reliable, structured tool use learned through reinforcement.
Our key contributions are as follows:
@@
    We introduce Random-Crypto, a novel procedurally generated cryptographic benchmark for LLM-agent training and evaluation.
@@
    We demonstrate that GRPO-based RL significantly improves tool-augmented reasoning in LLM agents.
@@
    We validate cross-domain generalization to picoCTF and AICrypto MCQ, revealing transferable, RL-acquired strategies for security-relevant problem solving across heterogeneous task formats.
We have released the benchmarks and training code used for this paper on GitHub. [^1] https://github.com/aielte-research/HackSynth-GRPO
@@ 2.1 CTF Challenges
Cryptographic challenges, in particular, revolve around exploiting weaknesses or logical flaws in encryption and hashing methods. These challenges commonly involve classical ciphers (e.g., Caesar, Vigenère, substitution), symmetric-key encryption (e.g., AES, DES), and public-key encryption schemes (e.g., RSA, ECC). Participants may also encounter tasks involving hashing algorithms, digital signatures, and custom cryptosystems, which often require creative approaches and deep theoretical understanding to solve.
@@ 2.2 LLM Agents in Cybersecurity
Recent advancements in LLMs have sparked significant interest in their application within cybersecurity, particularly in penetration testing and security assessments. Researchers have adapted Capture The Flag (CTF) challenges as standardized benchmarks [5, 6, 8, 7], to evaluate the cybersecurity capabilities of LLM-based agents systematically. Initial evaluations revealed that general-purpose LLMs struggled to effectively handle specialized cybersecurity tasks due to their inherent inability to execute code or interface with operating systems and external tools [13, 5, 6, 14].
In other domains, to bridge this interaction gap, newer architectures introduced tool augmentation, empowering LLM agents to invoke external tools or APIs. Frameworks like ReAct (Reason+Act) demonstrated substantial improvements by integrating logical reasoning with actionable tool use [15]. OpenAI’s ChatGPT further extended this approach through a plugin/function-calling API, enabling deterministic JSON tool calls [16]. Additionally, Anthropic’s Model Context Protocol (MCP) provides a structured interface for models to interact seamlessly with external data sources and tools, facilitating complex cybersecurity tasks that require real-time computational support [17, 18]. These tool-augmented LLM agents have significantly advanced the state-of-the-art. However, they have yet to be fully utilized in autonomous penetration testing.
@@ 2.3 Reinforcement Learning to Enhance LLM Reasoning
Recent work has shown that traditional RLHF, while effective in improving helpfulness and safety, often falls short in domains that require structured, multi-hop reasoning like math, programming, and cybersecurity. To address this, specialized algorithms have emerged, such as Group Relative Policy Optimization (GRPO) [10]. GRPO is a variant of PPO tailored for reasoning-intensive tasks: instead of relying on scalar preference labels, it compares multiple candidate outputs and assigns relative rewards within a group. This relative evaluation strategy encourages diversity and robustness in generated solutions, helping the model refine its internal decision-making process. GRPO has been successfully applied to train DeepSeek-R1 [11], achieving state-of-the-art performance on mathematical reasoning and competitive coding benchmarks. Its low computational overhead and compatibility with low-rank adaptation methods such as QLoRA [21] also make it an attractive choice for resource-constrained environments.
@@ 3.1 Random-Crypto Benchmark Dataset
We introduce the Random-Crypto benchmark, a programatically generated dataset created for training LLM-based agents with Reinforcement Learning. The benchmark was designed to allow generating near infinite variations of CTF challenges based on 50 different cryptographic schemes. Solving the challenges requires the ability to exploit vulnerabilities in cryptographic schemes and protocols. Each challenge requires recovering a hidden flag in the form flag{...} by leveraging a specific weakness.
@@
  • Hard: Complex tasks requiring advanced cryptographic knowledge, scripting capabilities, and potentially external online tools or frequency analysis methods (e.g., substitution cipher with frequency analysis or a vulnerable custom cryptosystem).
@@
Each challenge is generated by first selecting a sub-type and then randomizing parameters specific to the cryptographic scheme, such as cipher keys or substitution dictionaries. Subsequently, a random flag is encrypted using the chosen parameters. An LLM-generated narrative accompanies each ciphertext, providing context and essential information to ensure solvability (see Appendix B, Prompt 1). For instance, a Caesar cipher challenge randomly selects a shift value between 1 and 25, generates a random plaintext flag, encrypts it, and embeds it within an LLM-generated short narrative.
To verify the solvability of each challenge, the authors conducted manual validation; at least one instance per sub-type was solved. Parameter spaces were also reviewed to avoid trivial or unsolvable configurations. The dataset comprises 50 manually validated challenges used exclusively for testing, along with 5,000 automatically generated challenges designated for agent training.
@@ 3.2 Reinforcement Learning
We fine-tuned the Llama-3.1-8B-Instruct [12] model for 250 training steps using the Group Relative Policy Optimization (GRPO) algorithm. The training specifically leveraged a constrained, multi-step code execution interface with a Python execution server via Anthropic’s Model Context Protocol (MCP), allowing the model up to four sequential tool invocations per challenge. During training, we restricted the training problems to only come from the easy subset of all the challenges. The decision to restrict training to easy challenges was motivated by initial experiments showing that more difficult challenges negatively impacted the training process. Models tended to converge prematurely to local optima by prioritizing superficial reward types, i.e., outputs matching formatting or syntax rewards. The rarity of the accuracy reward in the case of these hard challenges caused the models not to pursue solving them.
During each training step, we generated eight candidate trajectories for four distinct challenges, resulting in a batch size of 32. Each candidate trajectory consisted of iterative reasoning followed by structured JSON-formatted tool calls. The model interaction cycle included:
@@
The composite reward function described in Table 1 guided the model’s learning process, designed to encourage correct and efficient problem-solving behavior. To address the challenge of the model prematurely optimizing towards superficial rewards (such as formatting or tool-calling correctness without genuine problem-solving), we implemented a deduction strategy. Specifically, we penalized scenarios where the model produced a fictitious or made-up output immediately following a tool call without actually sending it to the MCP server. Furthermore, we penalized the model when it produced a hallucinated flag as the answer, even before any reasoning or tool calls happened.
The total token generation was limited to 8192 per interaction sequence, and training was augmented using QLoRA [21] adapters on one A100 80GB GPU.
We explicitly instructed the model to strictly adhere to the interaction protocol with structured prompts (see Appendix B, Prompt 2). The available Python execution server exposed three functionalities:
  • execute_python: Executes provided Python code and returns output.
  • list_variables: Lists variables currently stored in memory.
@@
Tool Calling Reward The model produces a valid JSON tool call 0.2
Execution Reward The MCP server executes the code without error 0.5
3.3 Evaluation Setup
We evaluate both baseline and GRPO-trained agents on three benchmarks: (i) the Random-Crypto test split, consisting of 50 manually validated cryptographic challenges, (ii) the public picoCTF benchmark [6] containing 120 heterogeneous CTF problems across six categories, and (iii) the AICrypto MCQ benchmark [22], a 135-question multiple-choice dataset covering a broad spectrum of cryptography topics, including both theoretical and computational problems. All evaluations use the same tool-augmented prompting format described in Appendix B, with eight independent generations per task, and are scored using Pass@8 and Maj@8 metrics.
4 Results
This section presents the evaluation of five leading LLMs: Llama-4-Scout-17B-16E [23], Llama-3.1-70B [24], Llama-3.1-8B [12], GPT-4.1 [25], and o3 [26]—on the randomized cryptographic benchmark. Additionally, we analyze the impact of RL-based fine-tuning on model performance.
4.1 Benchmarking
We report two standard LLM evaluation metrics: Pass@8 is the proportion of tasks for which at least one out of eight independent model generations successfully solves the task, and Maj@8 is the proportion of tasks for which at least five out of eight independent model generations successfully solve the task. Table 2 summarizes the experimental outcomes, measured by Pass@8 and Maj@8 metrics. GPT-4.1 and o3 models consistently outperform Llama-family models, highlighting the advanced reasoning capabilities of these architectures.
@@ 4.1 Benchmarking
Interestingly, while supplementary hints typically improve performance for scenarios without tool access, this trend reverses when tools are available, especially for smaller models like Llama-3.1-8B. In these cases, hints often prompted models to attempt brute-force approaches. For example, challenges requiring enumeration or brute-forcing strategies (such as ROT13 decoding or poorly salted hashes) led smaller models to generate inefficient Python scripts. Such attempts often resulted in the models generating Python scripts that tried to enumerate exhaustive lists of possible solutions or unnecessary computations, causing memory exhaustion and crashes in the execution environment. Conversely, larger and more capable models, such as GPT-4.1 and Scout, demonstrated improved performance when hints were available alongside tools, as they effectively incorporated hints into structured, computationally efficient reasoning processes.
Differences in prompt engineering also affected model performance. Specifically, the prompts utilized in this experiment were optimized primarily for Llama-3.1-8B, inadvertently disadvantaging larger Llama-family models.
@@ 4.2 Reinforcement Learning
To investigate the impact of Reinforcement Learning (RL) on cryptographic task performance, we conducted three distinct training experiments using the Llama-3.1-8B model: (i) training without hints for 251 steps on easy challenges, (ii) training with hints for 251 steps, (iii) curriculum-based training, initially providing hints for 130 steps, followed by 121 steps without hints.
Initially, the model often generated valid Python code through tool calls but would simultaneously provide incorrect or imagined answers instead of executing the code. As a result, even when the generated scripts could produce the correct outputs, the model’s premature hallucinations led to the termination of the interaction.
Figure 1 illustrates the progression of average rewards throughout training, showing an increase from approximately 0.6 to 0.9 across all scenarios. The hinted training regime consistently achieved higher rewards due to the hints guiding the model to the appropriate solution. The exact breakdown of the different rewards gained during training is shown in Figure 2. Metrics tracking successful Python code execution (defined as code running without errors) improved notably from roughly 39% to approximately 56%, with most of this improvement happening in the first 30 steps of training. The adherence to the correct answer format saw an even more substantial improvement, rising from about 40% to nearly 80%. This improvement happened in the first 50 steps. However, properly formatted tool calls remained steady at approximately 95%. The reason behind this is that the prompt was constructed to direct the model to produce proper tool calls. The rate of successful challenge solutions went from 10% to 30%, and unlike the other rewards, this was gradual during the full training, with variance across training steps.
Figure 1: Reward gained during training. The bright lines mark the average, while the shaded lines mark the actual data points. One data point shows the average of all rewards given out in a training step to all benchmarks.
Figure 2: Reward types gained during training. The bright lines mark the average, while the shaded lines mark the actual data points. One data point shows the average of all rewards given out in a training step to all benchmarks. We can observe the biggest threefold improvement in the accuracy of the model, indicating successful challenge resolution.
Post-training results on the easy subset of the Random-Crypto benchmark (Table 3) reveal significant performance improvements. The curriculum and hinted models both gained a significant 0.80 improvement to their Pass@8 scores, which is better than GPT-4.1 and only slightly behind the o3 model. The training setup that included no hints achieved a slightly lower Pass@8 score. On the contrary, it produced a significantly higher Maj@8 score, indicating that this training setup produces a more consistent model. Training with hints could guide the model toward accurate solutions more frequently, even when the model couldn’t guess it by chance. However, this could also make the model rely on the guidance too much and lower the performance when no hint is given. Overall, results indicate that it is better to leave out hints during the training phase, as the increase in the Pass@8 score is not worth the consistency shown by the Maj@8 score.
Table 3: Pass@8 and Maj@8 results after GRPO training on the Random-Crypto.
After Training Improvement
@@
w/ Hints 0.90 $0.14$ +0.80 $0.12$
w/o Hints $0.88$ 0.24 $0.78$ +0.22
Curriculum 0.90 $0.14$ +0.80 $0.12$
4.3 Generalization
To examine whether the gains obtained on Random-Crypto translate beyond the synthetic distribution, we evaluate the GRPO-fine-tuned agent on the public picoCTF benchmark [6]. Unlike Random-Crypto, which focuses purely on cryptographic primitives, picoCTF spans six heterogeneous categories: General, Cryptography, Web, Forensics, Binary Exploitation, and Reverse Engineering, thereby probing transfer to a substantially broader problem surface. This choice is deliberate: success on picoCTF requires the agent to (i) recognize when a challenge can be reduced to stand-alone Python reasoning, and (ii) invoke the REPL tool robustly under noisier prompts and domain shifts.
Of the 120 picoCTF challenges, 95 embed auxiliary artifacts (e.g., ELF binaries, PCAP traces, PNGs) that our current agent cannot yet process efficiently. Each task is attempted eight times with temperature 0.7 and nucleus sampling ($p=0.95$), and we report Pass@8 and Maj@8 as before.
Table 4 shows that the GRPO model trained with hints or the curriculum training setup increases Pass@8 from 0.10 to 0.18. Also Maj@8 score was increased by 0.03 with the hinted training setup. Notably, the training setup without hints achieved the least significant improvement.
Table 4: RL training’s effect on performance on the external picoCTF benchmark.
After Training Improvement